Fermi National Laboratory
 

Accelerator Update for June 19, 2006 - June 20, 2006

 

Monday June 19

The day shift began with Operations monitoring store 4781 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) not stacking due to a LCW problem with HV102, with the Booster, Main Injector (MI), and Recycler conducting studies, with MiniBooNE off due to a Horn power supply ground fault, and with NuMI taking beam.

Pbar resumed stacking at 8:56 AM. Experts investigated why the stacking rate was low.

Water technicians searched the Meson cooling towers at MS2 for a LCW leak at 10:35 AM. They didn’t find one. They will search through the M01 – M05 enclosures.

A RF expert began an upgrade of all the MI RF stations at 10:38 AM. He was installing a tube power dissipation module on each station. He completed a portion of this work at 11:06 AM.

Booster experts began at 12:45 PM changing a kicker (MKS04) into a second notcher.

At 2:02 PM, Operations reported that Ecool would be left off except for when electron cooling was actually needed. The Pelletron temperature was getting too warm.

Water techs reported at 2:14 PM that they found a leaky valve on a Meson quadrupole, (MC5Q2).

Operations began preparing the Recycler at 3:26 PM to stash antiprotons.

At 3:59 PM, Operations reported that MiniBooNE was still off and would be until Wednesday.

The Recycler stash ended and Pbar resumed stacking at 5:12 PM.

The Recycler began preparing to stash at 10 PM. The stash ended and Pbar resumed stacking at 10:54 PM.

Tuesday June 20

The midnight shift began with Operations monitoring store 4781 and the stash, with Pbar stacking, with CDF and D0 using the colliding beam, with MiniBooNE off due to a Horn power supply ground fault, and with NuMI taking beam.

Around 3 AM, the Proton Source tripped off. Many devices alarmed, from power supplies to vacuum valves, but some of them hadn’t actually tripped off. Operators reset everything that needed resetting.

The Recycler began preparing to stash at 4:04 AM. The stash ended at 5:34 AM.

At 5:40 AM, the Pelletron tripped off and would not reset. Operations called in an expert. With the Pelletron off, the Recycler can’t cool the antiprotons before sending them to the TeV.

At 6:27 AM, a Linac quadrupole power supply tripped off. Operators tried several repairs with no success. Operations called in a Linac tech. Later in the morning, an expert reported that the magnet had a short and further testing was continuing.

A Recycler expert reported at 7:55 AM that the Pelletron was ready for the next shot.

The Plans for Tuesday

The plans for today are to repair Linac and establish a store. It looks like the D0 access will occur tomorrow, 6/21/06. The access will begin at 6:30 AM and last about sixteen hours.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/21/2006    email Fermilab